The Brisbane Manufacturing Company produces a single model of a CD player. Each player is sold for $204 with a resulting contribution margin of $70.

Brisbane's management is considering a change in its quality control system. Currently, Brisbane spends $42,000 a year to inspect the CD players. An average of 2,200 units turn out to be defective - 1,760 of them are detected in the inspection process and are repaired for $80. If a defective CD player is not identified in the inspection process, the customer who receives it is given a full refund of the purchase price.

The proposed quality control system involves the purchase of an x-ray machine for $190,000. The machine would last for five years and would have salvage value at that time of $22,000. Brisbane would also spend $450,000 immediately to train workers to better detect and repair defective units. Annual inspection costs would increase by $22,000. This new control system would reduce the number of defective units to 400 per year. 330 of these defective units would be detected and repaired at a cost of $43 per unit. Customers who still received defective players would be given a refund equal to 120% of the purchase price.

Questions 1 & 2
1. What is the Year 3 cash flow if Brisbane keeps using its current system?

2. What is the Year 3 cash flow if Brisbane replaces its current system?


Questions 3 & 4
[NOTE: When computing present values, use the present value factors from the tables on page 118 in the Coursepack]
3. Assuming a discount rate of 8%, what is the net present value if Brisbane keeps using its current system?

4. Assuming a discount rate of 8%, what is the net present value if Brisbane replaces its current system?
Answer #1

cash flow for current system

Total defective units-detected

=2200-1760

=440units

year 3 cash flow

Annual inspection cost -repairing cost *units repaired- refunded units

= -42000 -(80*1760)-440*204

=-42000-140800-89760

= -272560$

new system cash flow in year 3

annual inspection cost -repairing cost*units repaired - refunded units

= -[42000+22000increase]-330*43-[400-330]*204*120%

= -64000-14190-17136

= -95326

NPV

CURRENT SYSTEM

YEAR cash flow pv factor at 8% presnt value
1-5 -272560 3.9927 -1088250

new system

YEAR cash flow pv factor at 8% presnt value
1-5 -95326 3.9927 -380608[95326*3.9927]
0 -190000 1 -190000
0 -450000 1 -450000
5 22000 0.6806 14973
NET PRESENT VALUE .-1005635$

NEW SYSTEM IS BETTER AS IT REULTS IN LOWER CASH OUTFLOW